OpenStudy (anonymous):

For the function f(x)= (3x-3)/(2x+7), find... the horizontal asymptote... y=____ and the vertical asymptote... x=_____ **for horizontal, the exponents are equal right? so 3/2 is the horizontal asymptote? and for the vertical, i have to factor it out right? but i feel like it's not factorable? what do you think? :) thanks!!

jimthompson5910 (jim_thompson5910):

the horizontal asymptote is correct

OpenStudy (ybarrap):

you are correct on the horizontal

OpenStudy (anonymous):

ooh yay! :)

OpenStudy (anonymous):

hurray!! XD

jimthompson5910 (jim_thompson5910):

to find the vertical asymptote, set the denominator equal to zero and solve for x

OpenStudy (anonymous):

haha wooowoo :P except I'm lost on the vertical one lol so woowoo-woo=woo :P

OpenStudy (anonymous):

ohhh okay so 2x+7=0 -7 -7 2x=7 x=7/2 ?

OpenStudy (agent0smith):

To find the vertical, set the denominator equal to zero and solve.

jimthompson5910 (jim_thompson5910):

this is because you cannot divide by zero values that cause division by zero errors yield vertical asymptotes

jimthompson5910 (jim_thompson5910):

2x+7 = 0 doesn't lead to 2x = 7

OpenStudy (anonymous):

ohh okay :) oops... my bad! 2x= -7 x= -7/2 ?

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

haha okay :) wait so if it's factorable, there would be two vertical asymptotes?

OpenStudy (anonymous):

and if it's not, just set it equal to 0 and solve for x?

jimthompson5910 (jim_thompson5910):

yes, assuming nothing cancels

OpenStudy (anonymous):

okay :) great!! Thank you!! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

okay :) great!! Thank you!! :)